OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 19 feet, what is its lateral surface area? Use π = Answer 836 ft2 990 ft2 1672 ft2 1980 ft2

Directrix (directrix):

Area of the circular base = pi *r^2 154 = 3.14 r^2 r = 7.003 LA of cylinder = perimeter of base times height LA = 2 pi (7.003) (19) LA = 835.59 square feet or 836 square feet

OpenStudy (anonymous):

thanks